# FoundationStereo: Zero-Shot Stereo Matching

### v1

@@ 2 Related Work
[…]
Rendering Realism High Low High High High High High
Scenes 10 9 0 4 18 3 12
Layout Realism Medium Low Low High High Medium High
Stereo Pairs 1K 40K 200K 103K 306K 62K 1000K
Resolution $1024\times 436$ $960\times 540$ $1920\times 1080$ $960\times 540$ $640\times 480$ $960\times 540$ $1280\times 720$
@@
Camera Params Constant Constant Constant Constant Constant Constant Varying
Table 1: Synthetic datasets for training stereo algorithms (excluding test images with inaccessible ground truth). Indicates reduced diversity, caused by including many similar frames from video sequences. Our dataset includes varying intrinsics and baseline.
@@ 3.1 Monocular Foundation Model Adaptation
Fig. 3 visualizes the power of rich monocular prior that helps to reliably predict on ambiguous regions which is challenging to deal with by naive correspondence search along the epipolar line.
@@ 3.2 Attentive Hybrid Cost Filtering
Axial-Planar Convolution (APC) Filtering. An hourglass network consisting of 3D convolutions, with three down-sampling blocks and three up-sampling blocks with residual connections, is leveraged for cost volume filtering. While 3D convolutions of kernel size $3\times 3\times 3$ are commonly used for relatively small disparity sizes [66, 22, 7], we observe it struggles with larger disparities when applied to high resolution images, especially since the disparity dimension is expected to model the probability distribution for the initial disparity prediction. However, it is impractical to naively increase the kernel size, due to the intensive memory consumption. In fact, even when setting kernel size to $5\times 5\times 5$ we observe unmanageable memory usage on an 80 GB GPU. This drastically limits the model’s representation power when scaling up with large amount of training data. We thus develop “Axial-Planar Convolution” which decouples a single $3\times 3\times 3$ convolution into two separate convolutions: one over spatial dimensions (kernel size $K_{s}\times K_{s}\times 1$) and the other over disparity ($1\times 1\times K_{d}$), each followed by BatchNorm and ReLU. APC can be regarded as a 3D version of Separable Convolution [14] with the difference that we only separate the spatial and disparity dimensions without subdividing the channel into groups which sacrifices representation power. The disparity dimension is specially treated due to its uniquely encoded feature comparison within the cost volume. We use APC wherever possible in the hourglass network except for the down-sampling and up-sampling layers.
@@ 4.1 Implementation Details
We implement FoundationStereo in PyTorch. The foundation model is trained on a mixed dataset consisting of our proposed FSD, together with Scene Flow [40], Sintel [4], CREStereo [31], FallingThings [56], InStereo2K [1] and Virtual KITTI 2 [5]. We train FoundationStereo using AdamW optimizer [36] for 200K steps with a total batch size of 128 evenly distributed over 32 NVIDIA A100 GPUs. The learning rate starts at 1e-4 and decays by 0.1 at 0.8 of the entire training process. Images are randomly cropped to 320$\times$736 before feeding to the network. Data augmentations similar to [33] are performed. During training, 22 iterations are used in GRU updates. In the following, unless otherwise mentioned, we use the same foundation model for zero-shot inference using 32 refinement iterations.
@@ 4.2 Benchmark Datasets and Metric
Datasets. We consider five commonly used public datasets for evaluation: Scene Flow [40] is a synthetic dataset including three subsets: FlyingThings3D, Driving, and Monkaa. Middlebury [47] consists of indoor stereo image pairs with high-quality ground-truth disparity captured via structured light. ETH3D [48] provides grayscale stereo image pairs covering both indoor and outdoor scenarios. KITTI 2012 [18] and KITTI 2015 [41] datasets feature real-world driving scenes, where sparse ground-truth disparity maps are provided, which are derived from LIDAR sensors.
@@ 4.3 Zero-Shot Generalization Comparison
Ours (Scene Flow) 5.5 1.8 3.2 4.9
Selective-IGEV* [62] 7.5 3.4 3.2 4.5
Ours 1.2 1.4 1.9 2.2
@@ 4.5 Ablation Study
Table 7: Left: Ablation study of proposed network modules. Right: Ablation study of whether to use SAD dataset when training the foundation model described in Sec. 4.1. The choices adopted in our full model are highlighted in green.
@@ 5 Conclusion
We introduced FoundationStereo, a foundation model for stereo depth estimation that achieves strong zero-shot generalization across various domains without fine-tuning. We envision such a foundation model will facilitate broader adoption of stereo estimation models in practical applications. Despite its remarkable generalization, our method is not without limitations. First, our model is not yet optimized for efficiency, which takes 0.7s on image size of 375$\times$1242 on NVIDIA A100 GPU. Future work could explore adapting distillation and pruning techniques applied to other vision foundation models [82, 11]. Second, our dataset FSD includes a limited collection of transparent objects. Robustness could be further enhanced by augmenting with a larger diversity of fully transparent objects during training.
@@ 5.1 ETH3D Leaderboard
Figure 7: Examples scene models involving factory, hospital, wood attic, office, grocery store and warehouse. In the third column, we demonstrate an example of metallic material randomization being applied to augment scene diversity. The last column shows comparison of a warehouse between the real and our simulated digital twin in high fidelity.
5.2 More Ablation Study on Synthetic Data
We study the effectiveness of self-curation pipeline introduced in Sec. 3.5. When disabling the self-curation while keeping the same data size, the synthetic dataset involves ambiguous samples that confuse the learning process, leading to slight performance drop when evaluated on Middlebury [47] dataset.
@@ 5.2 More Ablation Study on Synthetic Data
5.3 More Details of Synthetic Data Generation
@@
Disparity Distribution. Fig. 8 shows the disparity distribution of our SAD dataset.
5.4 Acknowledgement
We would like to thank Gordon Grigor, Jack Zhang, Karsten Patzwaldt, Hammad Mazhar and other NVIDIA Isaac team members for their tremendous engineering support and valuable discussions.

### v3

@@ 2 Related Work
Simulator Blender Blender Blender Unreal Engine Unreal Engine Unreal Engine Unreal Engine Blender NVIDIA Omniverse
Rendering Realism High Low High High High High High High High
Scenes 10 9 0 4 18 3 8 47 12
Layout Realism Medium Low Low High High Medium High High High
Stereo Pairs 1K 40K 200K 103K 306K 62K 7.7K 6K 1000K
Resolution $1024\times 436$ $960\times 540$ $1920\times 1080$ $960\times 540$ $640\times 480$ $960\times 540$ $3840\times 2160$ $1920\times 1080$ $1280\times 720$
@@
Camera Params Constant Constant Constant Constant Constant Constant Constant Constant baseline, varying intrinsics Varying baseline and intrinsics
Table 1: Synthetic datasets for training stereo algorithms (excluding test images with inaccessible ground truth). Indicates reduced diversity, caused by including many similar frames from video sequences.
@@ 3.1 Monocular Foundation Model Adaptation
Fig. 3 visualizes the power of rich monocular prior that helps to reliably predict on ambiguous regions which is challenging to deal with by naive correspondence search along the epipolar line. Instead of using the raw monocular depth from DepthAnythingV2 which has scale ambiguity, we use its latent feature as geometric priors extracted from both stereo images and compared through cost filtering as described next.
@@ 3.2 Attentive Hybrid Cost Filtering
Axial-Planar Convolution (APC) Filtering. An hourglass network consisting of 3D convolutions, with three down-sampling blocks and three up-sampling blocks with residual connections, is leveraged for cost volume filtering [69, 1]. While 3D convolutions of kernel size $3\times 3\times 3$ are commonly used for relatively small disparity sizes [69, 23, 8], we observe it struggles with larger disparities when applied to high resolution images, especially since the disparity dimension is expected to model the probability distribution for the initial disparity prediction. However, it is impractical to naively increase the kernel size, due to the intensive memory consumption. In fact, even when setting kernel size to $5\times 5\times 5$ we observe unmanageable memory usage on an 80 GB GPU. This drastically limits the model’s representation power when scaling up with large amount of training data. We thus develop “Axial-Planar Convolution” which decouples a single $3\times 3\times 3$ convolution into two separate convolutions: one over spatial dimensions (kernel size $K_{s}\times K_{s}\times 1$) and the other over disparity ($1\times 1\times K_{d}$), each followed by BatchNorm and ReLU. APC can be regarded as a 3D version of Separable Convolution [15] with the difference that we only separate the spatial and disparity dimensions without subdividing the channel into groups which sacrifices representation power. The disparity dimension is specially treated due to its uniquely encoded feature comparison within the cost volume. We use APC wherever possible in the hourglass network except for the down-sampling and up-sampling layers.
@@ 4.1 Implementation Details
We implement FoundationStereo in PyTorch. The foundation model is trained on a mixed dataset consisting of our proposed FSD, together with Scene Flow [41], Sintel [5], CREStereo [32], FallingThings [59], InStereo2K [2] and Virtual KITTI 2 [6]. We train FoundationStereo using AdamW optimizer [37] for 200K steps with a total batch size of 128 evenly distributed over 32 NVIDIA A100 GPUs. The learning rate starts at 1e-4 and decays by 0.1 at 0.8 of the entire training process. Images are randomly cropped to 320$\times$736 before feeding to the network. Data augmentations similar to [34] are performed. During training, 22 iterations are used in GRU updates. In the following, unless otherwise mentioned, we use the same foundation model for zero-shot inference with 32 refinement iterations and 416 for maximum disparity.
@@ 4.2 Benchmark Datasets and Metric
Datasets. We consider five commonly used public datasets for evaluation: Scene Flow [41] is a synthetic dataset including three subsets: FlyingThings3D, Driving, and Monkaa. Middlebury [49] consists of indoor stereo image pairs with high-quality ground-truth disparity captured via structured light. Unless otherwise mentioned, evaluations are performed on half resolution and non-occluded regions. ETH3D [50] provides grayscale stereo image pairs covering both indoor and outdoor scenarios. KITTI 2012 [19] and KITTI 2015 [43] datasets feature real-world driving scenes, where sparse ground-truth disparity maps are provided, which are derived from LIDAR sensors.
@@ 4.3 Zero-Shot Generalization Comparison
[…]
@@ 4.4 In-Domain Comparison
In addition, our finetuned model also ranks 1st on the Middlebury leaderboard. See appendix for details.
@@ 4.5 Ablation Study
Table 7: Left: Ablation study of proposed network modules. Right: Ablation study of whether to use FSD dataset when training the foundation model described in Sec. 4.1. The choices adopted in our full model are highlighted in green.
@@ 5 Conclusion
We introduced FoundationStereo, a foundation model for stereo depth estimation that achieves strong zero-shot generalization across various domains without fine-tuning. We envision such a foundation model can facilitate broader adoption of stereo estimation models in practical applications. Despite its remarkable generalization, it has several limitations. First, our model is not yet optimized for efficiency, which takes 0.7s on image size of 375$\times$1242 on NVIDIA A100 GPU. Future work could explore adapting distillation and pruning techniques applied to other vision foundation models [85, 12]. Second, our dataset FSD includes a limited collection of transparent objects. Robustness could be further enhanced by augmenting with a larger diversity of fully transparent objects during training.
@@ 6 ETH3D Leaderboard
Figure 7: Examples scene models involving factory, hospital, wood attic, office, grocery store and warehouse. In the third column, we demonstrate an example of metallic material randomization being applied to augment scene diversity. The last column shows comparison of a warehouse between the real (bottom) and our simulated digital twin (top) in high fidelity.
7 Middlebury Leaderboard
At the time of submission, our fine-tuned model ranks 1st on the Middlebury leaderboard, significantly outperforming both published and unpublished works. The screenshot is shown in Fig. 8.
Figure 8: Middlebury leaderboard screenshot. Our fine-tuned foundation model (red box) ranks 1st at the time of submission.
Figure 9: Disparity distribution in our proposed FSD.
8 More Ablation Study on Synthetic Data
Effects of Self-Curation. We study the effectiveness of self-curation pipeline introduced in Sec. 3.5. When disabling the self-curation while keeping the same data size, the synthetic dataset involves ambiguous samples that confuse the learning process, leading to slight performance drop when evaluated on Middlebury [49] dataset.
@@ 8 More Ablation Study on Synthetic Data
Effects of FSD for Other Methods. We train representative works IGEV and Selective-IGEV on FSD only. As shown in the table below, for both methods, our proposed FSD effectively boosts the performance compared to the commonly used Scene Flow dataset.
[…]
IGEV FSD 7.8 3.5 3.2 4.7
Selective-IGEV Scene Flow 9.2 5.7 4.5 5.6
[…]
Table 9: Effects of FSD for other methods.
9 Results on Translucent Objects
We evaluate on Booster [46] (half resolution), which is a challenging dataset consisting of specular and transparent objects. We compare with the most competitive methods from Fig. 5 (main paper) in the zero-shot setting. The quantitative and qualitative results are shown below.
[…]
IGEV 30.8 22.3 19.0 22.7
[…]
10 More Results on Middlebury Dataset
We compare with competitive methods that released their public weights in zero-shot on Middlebury, shown in below table. Since NMRF [21] did not report their evaluated Middlebury resolution, we rerun their released weights on all resolutions. At full resolution, maximum disparity 320 is used for FoundationStereo. Across all resolutions, ours significantly outperforms baselines. We also report the peak memory usage and running time averaged across the dataset on the same hardware, particularly single GPU 3090. On half and quarter resolutions, our peak memory occurs at STA module. On full resolution, it occurs at DT module. Despite the speed limitation which is not the focus when developing this work, ours can successfully run on a desktop GPU. Pruning or distillation remains an interesting future work to improve speed and memory footprint.
[…]
IGEV[33] 13.1 6.3 2.06 8.8 1.6 0.53 6.4 0.5 0.18
IGEV++[66] 12.7 13.1 2.12 7.8 3.4 0.50 6.3 0.9 0.15
NMRF[20] 35.3 8.1 0.95 10.9 1.8 0.20 5.0 0.5 0.05
[…]
Table 10: Results on varying resolutions in Middlebury.
11 More Details of Synthetic Data Generation
@@
Disparity Distribution. Fig. 9 shows the disparity distribution of our FSD dataset.
12 Acknowledgement
We would like to thank Gordon Grigor, Jack Zhang, Xutong Ren, Karsten Patzwaldt, Hammad Mazhar and other NVIDIA Isaac team members for their tremendous engineering support and valuable discussions.